OpenStudy (anonymous):

find the area of the region bounded by the graphs of f(x)=x^2+2 and g(x)=4-x

OpenStudy (zpupster):

draw a graph set them equal to each other, so we can see where they intersect x^2+2 = 4-x x^2 +x -2 = 0 (x+2)(x-1) = 0 To find the area, we must integrate the difference between the upper function and lower function sub the values